Hi, Can you please me solve this problem thank you very much.
The question is:
A building hold 80 apartments. Une bedroom apartment cost $850/month. A two bedroom apartment cost 1050/month. The location itself cost 78 400/month. How many one bedrooms and two bedrooms are in this building?
Thank you again.
-kim
I am asking this question again because i wrote in the wrong amount instead of 78 200 , the actual amount is 78 400
thank you

Answer by wgomero(28)

resolution:
let x = 1 bedroom and y = 2 bedroom
1) x + y = 80;
2) 850x + 1050y = 78400
let solve equation 1 by y; y = 80 - x then plug in equation 2
850x + 1050(80 -x) = 78400
850x + 84000 - 1050x = 78400
-200x = 78400 - 84000
-200x = - 5600
x = 28
now find y in equation #1: 28 + y = 80; y = 80 - 28; y = 52
one bedroom = 28; two bedroom = 52.